rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#15122 — bhs2-16ab-n56

Attached to Project— Network
Planned Work
Beauharnois   → BHS-4
CLOSED
100%
Intervention date between 26 of October 08:00 AM and 10:00 UTC

We have CRC on one of the VPC links of this pair of n56, we will clean the optical fiber and replace the SFP modules if necessary.
No impact to predict the link is redundant.

Thank you for your understanding.
Date:  Thursday, 26 October 2017, 08:28AM
Reason for closing:  Done
Additional comments about closing:  Fiber cleanned no more CRC.
No impact.
Comment by OVH - Thursday, 26 October 2017, 08:20AM

We are currently cleaning the fiber.
The vPC interface Ethernet 1/47 is currently shutted. No impact a vPC link is still up.